OpenStudy (anonymous):

the region enclosed by the graph y=x^2, the line x=2,and the x-axis is rearopund the y-axis. what the volume of the solid generated?

OpenStudy (anonymous):

Here is my solution: \[\int\limits_{0}^{2}(2\pi)(x)(x^2)\] then, then that is \[(2\pi)\int\limits_{0}^{2}(x^4)/4\] then you receive 2pi*(4-0) and get \[8\pi\] someone should check this as this is just what I got. Hope it helps

OpenStudy (nikvist):

\[V=\pi\left(\int\limits_0^4 2^2dy-\int\limits_0^4(\sqrt{y})^2dy\right)=\pi\left(4\int\limits_0^4 dy-\int\limits_0^4 ydy\right)=\]\[=\pi\left(16-8\right)=8\pi\]